OpenStudy (anonymous):

Find all solutions in the interval [0, 2π). 2 sin^2 x = sin x

OpenStudy (mathmale):

I'd suggest you re-write the given equation first as:\[2 \sin^2 x = \sin x\] and next as\[2 \sin^2 x-\sin x=0\] and lastly as\[2 (\sin x)^2-\sin x=0\]

OpenStudy (mathmale):

Try factoring that. Just factor out sin x. Can you now find the two solutions (values of sin x)? Once you have those, solve for the angles x on the interval [0, 2Pi].

OpenStudy (anonymous):

Umm I'm going t otry to factor it out first

OpenStudy (anonymous):

sinx (sinx -1) ? I get a little confused wit factoring and sin's

OpenStudy (anonymous):

@mathmale

OpenStudy (mathmale):

OK. Let's take a different approach to make this problem look more familiar.

OpenStudy (anonymous):

alrighty

OpenStudy (mathmale):

We start with \[2 (\sin x)^2-\sin x=0\]and substitute w for sin x:

OpenStudy (mathmale):

\[2w^2-w=0\]

OpenStudy (anonymous):

2(w)^2 -w=0

OpenStudy (mathmale):

We factor that and get what factors?

OpenStudy (anonymous):

Since there isn't three "number"s how do I work it out ? Because I know the trick that you find two digits that when added equal the middle number

OpenStudy (anonymous):

w(2w-1)?

OpenStudy (anonymous):

@mathmale

OpenStudy (mathmale):

Yes. then either w = 0 or (2w-1)=0. solve the 2nd equation for w, please.

OpenStudy (anonymous):

=1/2

OpenStudy (anonymous):

2w+1=0 too?

OpenStudy (anonymous):

=-1/2

OpenStudy (mathmale):

Good. Now remember that we let w=sin x a few minutes ago. If w= 0, that also means that sin x = 0. At which values of x within the interval [0,2Pi] is sin x = 0?

OpenStudy (anonymous):

pi/2 and 3pi/2

OpenStudy (mathmale):

Also, 2w-1=0, or w = +1/2. this means that sin x = 1/2. There are two angles within the interval [0,2Pi] at which the sine function = 1/2. What are those 2 angles? You should end up with 5 solutions (3 of them come from sin x = 0 and 2 from sin x =+1/2). I need to get off the 'Net now, but if you're not satisfied that you 've solved this problem completely, leave a note for me, and then someone else or I will respond, later.

OpenStudy (mathmale):

I think you mean cos x = 0. the solutions to that are Pi/2 and 3Pi/2. But we were working with sin x = 0. What are the solutions to that? (there are 3 solutions on the interval [0,2Pi].)

OpenStudy (anonymous):

I was looking at the completed unit circle how do you know if it is cos or sin

OpenStudy (anonymous):

 Here are the points on a unit circle. If you want to find when sinx=0 then look where y is 0